Could you Please help me with this problem. Thanks, in advance.
The average number of seeds in a package of cucumber seeds is 25. The number of seeds in the package can vary by three. what are the maximum and minimum number of seeds that could be in a package.
Thank You,
Raj

Answer by stanbon(48546)

The average number of seeds in a package of cucumber seeds is 25. The number of seeds in the package can vary by three. what are the maximum and minimum number of seeds that could be in a package.
-------------
mimimum = 25-3 = 22
maximum = 25+3 = 28
=================
Cheers,
stan H.